OpenStudy (anonymous):

Let f(x) = x^2 – 16. Find f^–1(x).

OpenStudy (anonymous):

A) +_ 4 sqrt x B) +_ sqrt x + 16 C)x^2/16 D) 1/x^2-16

OpenStudy (anonymous):

let y = x^2 - 16. switch the positions of the x's and y's to get x = y^ - 16. then solve for y. x + 16 = y^2 sqrt(x+16) = y

OpenStudy (anonymous):

B, then